# Routine haematological markers can predict and discriminate health status and biological age even from noisy sources

### v9

@@ 3.4 UK Biobank Subset Validation
We first filtered the Rest (‘healthy’ control) cohort by removing all individuals whose third-time point CBC measurements deviated from established normal physiological ranges. This reduced the noise and the cohort size from $n=2312$ to $n=1741$. For disease groups (Cancer, CVD, Infection), we kept only those individuals who exhibited at least four CBC values that fall outside the normal range. This filtering reduced the initial cohort sizes (Cancer: 666, CVD: 632, Infection: 396) to 438, 417, and 367 respectively. Pairwise comparisons of composite NIS scores between these filtered groups are summarised in Table 12, and box plots of scores are presented in Fig. 7. The Shapiro–Wilk tests yielded p-values << 0.05 for all groups, thus we used the Kruskal–Wallis test followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons, as shown in Table 12.
@@ 4 Conclusions
The indexes/scores introduced here are not intended to be used as a diagnostic tool alone, but are intended as a warning and flagging system capable of first triaging and general blood health screening. We have shown that the index captures, in a single numerical value and a colour-coded scheme, the abnormal nature of blood results in terms of deviation from healthy reference values (absolute or adaptable). The index is sensitive to out-of-range values and increases its value and changes its colour as a function of how removed values are from lower and upper bounds of reference values according to the number of standard deviations from the medians. Although it has here been proven to be sensitive enough, it is not specific enough to quantify diseases or conditions.
Another aspect of this retrospective study is that it shows that while inter-person stability over longitudinal data points may be important [32, 26, 27], useful single snapshot risk assessment scores are possible and of practical use according to their Cohen’s d values.
@@
As this research did not involve human subjects and used only publicly available anonymised data, it was exempted from consideration by an institutional review board and from ethical approvals.
@@ 4.3 Linearisation and Separability
Figure 11 shows how male and female NHANES data can be distinguished by the scores using Red Blood Cell (RBC) counts as a health status predictor. The normal versus abnormal RBC ranges from the NHS were used to label the person as healthy or unhealthy. The Pearson $R^{2}$ value for RBC-dependent health status was found to be 0.176, for both sexes considered together. The Pearson $R^{2}$ values were 0.203 and 0.455, for male and female RBC counts considered separately, when comparing healthy and unhealthy groups.
@@
When healthy and unhealthy patients were separated using a subset of four CBC/FBC analyte measures, namely, MCH, HCT, neutrophil, and lymphocyte counts, as shown in Figure 13, we obtained the second best $R^{2}$ value of 0.075, for both sexes. The $R^{2}$ values were 0.050 and 0.123, for the male and female groups, respectively. We found that neutrophil count alone, as a single clinical parameter, delivered a performance nearly equal to that of the composite measure of these four analyte parameters, with an $R^{2}$ of 0.072 when including both sexes. The inclusion of other white blood cell (WBC) parameters, such as monocyte count, eosinophil, and basophil, did not outperform these correlates. We observed that different subsets of a CBC/FBC by health status separated the data. All reported measures were statistically significant, with a two-way ANOVA P<0.0001<0.0001.

### v10

@@ 3.4 UK Biobank Subset Validation
We first filtered the Rest (‘healthy’ control) cohort by removing all individuals whose third-time point CBC measurements deviated from established normal physiological ranges. This reduced the noise and the cohort size from $n=2312$ to $n=1741$. For disease groups (Cancer, CVD, Infection), we kept only those individuals who exhibited at least four CBC values that fall outside the normal range. This filtering reduced the initial cohort sizes (Cancer: 666, CVD: 632, Infection: 396) to 438, 417, and 367 respectively. Pairwise comparisons of composite NIS scores between these filtered groups are summarised in Table 13, and box plots of scores are presented in Fig. 7. The Shapiro–Wilk tests on the residuals yielded p-values << 0.05 for all groups, thus we used the Kruskal–Wallis test followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons, as shown in Table 13.
@@ 4 Conclusions
The indexes/scores introduced here are not intended to be used as a diagnostic tool alone, but are intended as a warning and flagging system capable of first triaging and general blood health screening. We fully concur that abnormal blood counts naturally warrant medical evaluation. Our index is designed as an adjunctive computational triage tool that quantifies deviations in a standardized manner, complementing clinical decision-making rather than replacing expert clinical judgment. We have shown that the index captures, in a single numerical value and a colour-coded scheme, the abnormal nature of blood results in terms of deviation from healthy reference values (absolute or adaptable). The index is sensitive to out-of-range values and increases its value and changes its colour as a function of how removed values are from lower and upper bounds of reference values according to the number of standard deviations from the medians. Although it has here been proven to be sensitive enough, it is not specific enough to quantify diseases or conditions.
Another aspect of this retrospective study is that it shows that while inter-person stability over longitudinal data points may be important [33, 27, 28], useful single snapshot risk assessment scores are possible and of practical use when to their Cohen’s d values are statistically significant like in this case.
@@
We further validated the NIS and NCCIS immune scores on a UK Biobank subset spanning diverse disease groups, including cancers, infections (bacterial, viral, fungal), and cardiovascular diseases (e.g., stroke, heart attack). As shown in Figure 7, the scores demonstrated statistically significant discriminatory power in separating healthy individuals from each disease group.
@@
From an implementation standpoint, these indices could be embedded into electronic health record (EHR) systems as automated warning or longitudinal monitoring modules, providing clinicians with real-time quantitative feedback on patient trajectories. Similar to widely used risk assessment scores, the proposed metric condenses deviations in blood test parameters into a single, symbolically interpretable, and directly comparable value for applications such as triage, where the alternative would be a first-come, first-served or random-order approach.
@@
All methods were performed in accordance with the relevant guidelines and regulations as outlined by the institutional policies. As this research did not involve human subjects and used only publicly available anonymised data, it was exempted from consideration by an institutional review board and from ethical approvals. As the analysis was strictly computational, no further institutional approval was required.
@@ 4.4 Linearisation and Separability
When healthy and unhealthy patients were separated using a subset of four CBC/FBC analyte measures, namely, MCH, HCT, neutrophil, and lymphocyte counts, we obtained the second best $R^{2}$ value of 0.075, for both sexes. The $R^{2}$ values were 0.050 and 0.123, for the male and female groups, respectively. We found that neutrophil count alone, as a single clinical parameter, delivered a performance nearly equal to that of the composite measure of these four analyte parameters, with an $R^{2}$ of 0.072 when including both sexes. The inclusion of other white blood cell (WBC) parameters, such as monocyte count, eosinophil, and basophil, did not outperform these correlates. We observed that different subsets of a CBC/FBC by health status separated the data. All reported measures were statistically significant, with a two-way ANOVA P<0.0001<0.0001. The weak correlation analysis suggests limited predictive power of CBC analytes alone for distinguishing healthy from unhealthy individuals, particularly when applied to heterogeneous populations with broadly defined health status using routine blood measures. Therefore, these findings highlight the rationale for proposing the immune score as a diagnostic tool, as it offers significantly stronger predictive power than individual analyte counts, repurposing routine, accessible blood tests into a more informative clinical measure.